OpenStudy (anonymous):

write an equation of a line that is perpendicular with the equation -2x+7y=14: @AriPotta

OpenStudy (anonymous):

this is new to me

OpenStudy (aripotta):

ok first you wanna put that in slope-intercept form

OpenStudy (aripotta):

just like the other question

OpenStudy (anonymous):

y = 2/7x + 2

OpenStudy (aripotta):

oh it's exactly the same equation lol

OpenStudy (anonymous):

lol yea

OpenStudy (aripotta):

so perpendicular lines have a negative reciprocal slope to each other. so what's this line's slope? y = 2/7x + 2?

OpenStudy (anonymous):

2/7

OpenStudy (aripotta):

yes and negative reciprocal means that's gonna be negative and flipped. idk how else to explain that lol. the original line has a slope of 2/7 and the perpendicular line will have a slope of -7/2

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

so any equation, so long as it has a slope of -7/2 will be ok for this question

OpenStudy (aripotta):

like you could put in y = -7/2x + 2 and it should work

OpenStudy (anonymous):

thank u

OpenStudy (aripotta):

no problemo :)